OpenStudy (anonymous):

find the value of the trigonometric function to four decimal places. sec 8

OpenStudy (phi):

is the 8 in degrees or radians? sec(x)= 1/cos(x) if you know what units 8 is in, use a calculator to find 1/cos(8) in the correct mode (radians or degrees)

OpenStudy (anonymous):

\[\large 5\pi /2<8<6\pi /2\]

OpenStudy (anonymous):

thank you